I can normally do complete the square problems but this one's thrown me for a loop be of the fact that it has a fraction.
What is the last term for w^2 - 7/3w + __

Answer by stanbon(75887)

What is the last term for w^2 - 7/3w + __
----------
Rule: Multiply the coefficient of the 1st degree term by (1/2) and square
to get the 3rd term of the trinomial.
------------
(1/2)(-7/3) = (-7/6)^2 = 49/36
===============
w^2 -(7/3)w + 49/36 is a perfect trinomial square.
=============
Cheers,
Stan H.